OpenStudy (anonymous):

what is the square root of 108 in simplified radical form?

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

ok so sqrt108. we need to find a biggest poosible perfect square number by which we can divide 108. this number is 36. 36=6*6. then sqrt(36*3)=sqrt(6*6*3)=\[6\times \sqrt{3}\]

OpenStudy (anonymous):

can I ask another question

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

yes but post it as a new post and also medla me